# Applications of an identity of Batır

### v1

@@ 2 General results
$\sum_{k=1}^{n}\frac{1}{k(k+p)}=\frac{1}{p}\sum_{k=1}^{n}\left(\frac{1}{k}-\frac{1}{k+p}\right)=\frac{1}{p}\left(H_{n}+H_{p}-H_{n+p}\right).$
@@ 3 Miscellaneous results involving binomial coefficients
$a_{k}=(-1)^{k}\binom{p}{n-k}\binom{k}{m}$
in (15), since ([12, Corollary 5])
$\sum_{k=1}^{n}(-1)^{k}\binom{p}{n-k}\binom{k}{m}H_{k}=(-1)^{n}\binom{{n-p}}{{m-p}}\left(H_{n-p}-H_{m-p}+H_{m}\right).$
@@
$\sum_{k=0}^{n}{\binom{n}{k}^{2}H_{k}}=\binom{{2n}}{n}\left(2H_{n}-H_{2n}\right).$
@@
$\sum_{k=0}^{n}{\binom{n}{k}\binom{2n}{k}H_{k}}=\binom{3n}{n}H_{n}-\sum_{k=1}^{n}\frac{1}{k}\binom{3n-k}{n-k}$.
@@
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}^{3}}=\begin{cases}(-1)^{n/2}\binom{{n}}{{n/2}}\binom{{3n/2}}{n},&\text{if $n$ is even;}\\ 0,&\text{if $n$ is odd.}\end{cases}$
@@
$\sum_{k=1}^{n}\binom{n+s}{k}(-1)^{k}H_{k}=(-1)^{n}\binom{n+s-1}{n}H_{n}+\frac{(-1)^{n}}{n+s}\binom{n+s-1}{n}-\frac{1}{n+s}.$
@@
$\sum_{k=1}^{n}\binom{s}{k}(-1)^{k}sH_{k}=s(-1)^{n}\binom{s-1}{n}H_{n}+(-1)^{n}\binom{s-1}{n}-\frac{1}{s}.$
@@
$\binom{1/2}{r}=(-1)^{r+1}\frac{2^{-2r}}{2r-1}\binom{2r}{r},\quad\binom{-1/2}{r}=(-1)^{r}2^{-2r}\binom{2r}{r}$,
@@
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}^{-1}H_{k}}=\left({n+1}\right)\left({\frac{{H_{n+1}}}{{n+2}}-\frac{{(-1)^{n}+1}}{{\left({n+2}\right)^{2}}}}\right).$
@@
$\binom{{k+q}}{p}=\binom{{k+q+1}}{{p+1}}-\binom{{k+q}}{{p+1}}$.
@@
$\binom{r}{s}=\frac{{\Gamma(r+1)}}{{\Gamma(s+1)\Gamma(r-s+1)}}.$
@@
$\binom{n+1/2}{j}=\frac{\binom{2n+1}{2j}\binom{2j}{j}}{\binom{n}{j}}2^{-2j}\quad\text{and}\quad\binom{n}{n+1/2}=\frac{1}{\pi}\frac{2^{2n+2}}{n+1}\binom{2(n+1)}{n+1}^{-1}.$
@@ 4 Applications to binomial transforms
$\sigma_{n}=\sum_{k=0}^{n}\binom{n}{k}(-1)^{k}s_{k}\qquad\iff\qquad s_{n}=\sum_{k=0}^{n}\binom{n}{k}(-1)^{k}\sigma_{k}.$
@@
$s_{n}=x^{n}H_{n}\quad\text{and}\quad\sigma_{n}=(1-x)^{n}H_{n}-\sum_{k=1}^{n}\frac{(1-x)^{n-k}}{k},$
@@
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\frac{{2k+1}}{{k+1}}C_{k}\left({H_{k+1}-O_{k+1}}\right)}=-2^{-2n}\frac{{2n+1}}{{n+1}}C_{n}\left({H_{n+1}-O_{n+1}}\right)$
@@
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k+1}}}{{k+1}}}=(-1)^{n+1}\frac{{B_{n+1}}}{{n+1}}-\frac{1}{{n+1}}$
@@
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{2k}}{k}}=(-1)^{n}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{k}\binom{{n-k}}{k}}$.

### v2

@@ 2 General results
$\sum_{k=1}^{n}\frac{1}{k(k+p)}=\frac{1}{p}\sum_{k=1}^{n}\left(\frac{1}{k}-\frac{1}{k+p}\right)=\frac{1}{p}\left(H_{n}+H_{p}-H_{n+p}\right).$ (32)
@@ 3 Miscellaneous results involving binomial coefficients
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{\sum_{j=\left\lfloor{n/2}\right\rfloor}^{k-1}{\frac{1}{{n-j}}\binom{{j}}{{n-j}}}}=F_{n+1}-1.$ (43)
Proof.
Use $a_{k}=\binom{n-k}{k}$ in Theorem 1, noting that [8, Equation (6.130), p.303]:
$\sum_{k=0}^{n}{\binom{n-k}{k}}=F_{n+1}.$ (44)
Proposition 14.
@@
$a_{k}=(-1)^{k}\binom{p}{n-k}\binom{k}{m}$ (46)
in (16), since ([13, Corollary 5])
$\sum_{k=1}^{n}(-1)^{k}\binom{p}{n-k}\binom{k}{m}H_{k}=(-1)^{n}\binom{{n-p}}{{m-p}}\left(H_{n-p}-H_{m-p}+H_{m}\right).$ (47)
@@
$\sum_{k=0}^{n}{\binom{n}{k}^{2}H_{k}}=\binom{{2n}}{n}\left(2H_{n}-H_{2n}\right).$ (55)
@@
$\sum_{k=0}^{n}{\binom{n}{k}\binom{2n}{k}H_{k}}=\binom{3n}{n}H_{n}-\sum_{k=1}^{n}\frac{1}{k}\binom{3n-k}{n-k}$. (56)
@@
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}^{3}}=\begin{cases}(-1)^{n/2}\binom{{n}}{{n/2}}\binom{{3n/2}}{n},&\text{if $n$ is even;}\\ 0,&\text{if $n$ is odd.}\end{cases}$ (58)
@@
$\sum_{k=1}^{n}\binom{n+s}{k}(-1)^{k}H_{k}=(-1)^{n}\binom{n+s-1}{n}H_{n}+\frac{(-1)^{n}}{n+s}\binom{n+s-1}{n}-\frac{1}{n+s}.$ (63)
@@
$\sum_{k=1}^{n}\binom{s}{k}(-1)^{k}sH_{k}=s(-1)^{n}\binom{s-1}{n}H_{n}+(-1)^{n}\binom{s-1}{n}-\frac{1}{s}.$ (64)
@@
$\binom{1/2}{r}=(-1)^{r+1}\frac{2^{-2r}}{2r-1}\binom{2r}{r},\quad\binom{-1/2}{r}=(-1)^{r}2^{-2r}\binom{2r}{r}$, (67)
@@
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}^{-1}H_{k}}=\left({n+1}\right)\left({\frac{{H_{n+1}}}{{n+2}}-\frac{{(-1)^{n}+1}}{{\left({n+2}\right)^{2}}}}\right).$ (70)
@@
$\binom{{k+q}}{p}=\binom{{k+q+1}}{{p+1}}-\binom{{k+q}}{{p+1}}$. (73)
@@
$\binom{r}{s}=\frac{{\Gamma(r+1)}}{{\Gamma(s+1)\Gamma(r-s+1)}}.$ (76)
@@
$\binom{n+1/2}{j}=\frac{\binom{2n+1}{2j}\binom{2j}{j}}{\binom{n}{j}}2^{-2j}\quad\text{and}\quad\binom{n}{n+1/2}=\frac{1}{\pi}\frac{2^{2n+2}}{n+1}\binom{2(n+1)}{n+1}^{-1}.$ (85)
@@ 4 Applications to binomial transforms
$\sigma_{n}=\sum_{k=0}^{n}\binom{n}{k}(-1)^{k}s_{k}\qquad\iff\qquad s_{n}=\sum_{k=0}^{n}\binom{n}{k}(-1)^{k}\sigma_{k}.$ (86)
@@
$s_{n}=x^{n}H_{n}\quad\text{and}\quad\sigma_{n}=(1-x)^{n}H_{n}-\sum_{k=1}^{n}\frac{(1-x)^{n-k}}{k},$ (97)
@@
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\frac{{2k+1}}{{k+1}}C_{k}\left({H_{k+1}-O_{k+1}}\right)}=-2^{-2n}\frac{{2n+1}}{{n+1}}C_{n}\left({H_{n+1}-O_{n+1}}\right)$ (100)
@@
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k+1}}}{{k+1}}}=(-1)^{n+1}\frac{{B_{n+1}}}{{n+1}}-\frac{1}{{n+1}}$ (102)
@@
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{2k}}{k}}=(-1)^{n}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{k}\binom{{n-k}}{k}}$. (104)
@@
Proposition 33.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{{(-1)^{j}}}{{n-j+1}}\binom{{n}}{j}C_{k-j+1}}}=\frac{1}{{n+1}}\left({\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}C_{k}}-(-1)^{n}}\right).$ (105)
Proof.
Use the result [1, Prop. 6]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}C_{k+1}}=(-1)^{n}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}C_{k}}$ (106)
in (94). ∎